# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 8. Conclusion
This paper contributes a comprehensive overview of ILFPPM, presenting a foundational understanding of privacy preservation in indoor location fingerprinting. By addressing various dimensions, including definitions, applications, vulnerabilities, and protection mechanisms, we aim to provide more insights into the complexities of this dynamic domain. We introduce all possible origins of privacy leakages in indoor location fingerprinting. The novel categorization of adversary and attack models, along with the compilation of datasets and metrics, serves as a valuable resource for next empirical investigations. Through a detailed exploration of challenges, we point out the future research directions and opportunities in the field of privacy-preserving mechanisms for indoor location fingerprinting. With this extensive survey, we aim to establish a robust groundwork for forthcoming studies in this area.

### v2

@@ 8 Conclusion
This paper contributes a comprehensive overview of indoor location fingerprinting privacy-preserving mechanisms (ILFPPM), presenting a foundational understanding of privacy preservation in indoor location fingerprinting (ILF) systems. By addressing various privacy dimensions in such systems, including definitions, applications, vulnerabilities, attacks, metrics, datasets, and protection mechanisms, we aim to provide more insights into the complexities of this dynamic domain. We introduce all possible origins of privacy leakages in ILF systems. The novel categorization of adversary and attack models, along with the compilation of datasets and metrics, serves as a valuable resource for the next empirical investigations. Through a detailed exploration of challenges, we point out future research directions and opportunities in the field of privacy-preserving mechanisms for ILF systems. With this extensive survey, we aim to establish a robust groundwork for forthcoming studies in this area.